# Configurations, Tessellations and Tone Networks

### v4

@@ Abstract
The Eulerian tonnetz, which associates three minor chords to each major chord and three major chords to each minor chord, can be represented by a bipartite graph with twelve white vertices denoting major chords and twelve black vertices denoting minor chords. This so-called Levi graph determines a configuration of twelve points and twelve lines in $\mathbb{R}^{2}$ with the property that three points lie on each line and three lines pass through each point. Interesting features of the tonnetz, such as the existence of the four hexatonic hexacycles and the three octatonic octacycles, crucial for the understanding of nineteenth-century harmony and voice leading, can be read off rather directly as properties of this $\{12_{3}\}$ and its Levi graph. Analogous tone networks together with their associated Levi graphs and configurations can be constructed for pentatonic music and twelve-tone music, offering the promise of new methods of composition. When the constraints of the Eulerian tonnetz are relaxed so as to allow movements between major and minor triads with variations at exactly two tones, the resulting bipartite graph has two components, each of which generates a tessellation of the plane, of a type known to Kepler, based on hexagons, squares and dodecagons. When the same combinatorial idea is applied to tetrachords of the Tristan genus (dominant sevenths and minor sixths) the cycles of the resulting bipartite graph are sufficiently ample in girth to ensure the existence of a second geometrical configuration of type $\{12_{3}\}$, distinct from the Eulerian tonnetz as an incidence geometry, which can be used as the basis for a new approach to the analysis of the music of Chopin, Wagner, Tchaikovsky, Brahms and their contemporaries.
@@ I Introduction
The well-known Eulerian tonnetz, which consists of a network of relations between the conventional major and minor chords of Western music, was originally introduced, in part form, by Leonhard Euler (1739) and later developed in works of Gottfried Weber (1824), Ernst Naumann (1858), Arthur von Oettingen (1866) and Hugo Riemann (1902).  [^1] We treat the word ‘tonnetz’, spelled with a lower case ‘t’, and its plural ‘tonnetze’ as words of English.   Investigation of the tonnetz from a modern standpoint in the setting of equal temperament was carried forward in works of Waller (1978), Lewin (1982, 1987), Hyer (1989), Cohn (1996, 1997, 1998), Douthett & Steinbach (1998) and many others, attracting much attention.
@@
Figure 1 shows a representation of the Eulerian tonnetz as a tessellation (or periodic tiling) of the plane by regular hexagons. This figure, viewed as an infinite graph, is familiar to music theorists – see, for example, Douthett & Steinbach (1998) at p. 248 or Tymoczko (2011), at p. 413. But the tessellation under consideration in Figure 1 is more than merely a graph: it is a geometrical object. We propose here to introduce another truly geometrical (not simply graph-theoretic or lattice-theoretic) representation of the tonnetz – specifically, we show that the tonnetz can be visualized as a configuration of twelve lines and twelve points in the Euclidean plane, with the property that three of the points lie on each line and three of the lines pass through each point. The construction proceeds in three steps. First, we interpret the tonnetz as a regular bipartite graph of degree three and girth six with the two types of vertices corresponding to major and minor triads, respectively, and the edges encoding adjacency within the tonnetz. This allows for easy representation and analysis of the various cyclic chord progressions that arise naturally in the tonnetz. We then go on to realize this Levi graph as an incidence geometry consisting of twelve points (major triads) and twelve lines (minor triads) in $\mathbb{R}^{2}$, in the form of a so-called configuration $\{12_{3}\}$, leading to interesting visualizations. Finally, we identify this configuration as the self-dual $\{12_{3}\}$ of type D222 described by Daublebsky von Sterneck (1895).
@@
The tonnetz is widely regarded as essential and indeed fundamental to the modern understanding of music based on triadic harmonies – so it is both unexpected and important to see such a connection between the tonnetz and a basic construction in Euclidean geometry. But there are challenges as well. Can the idea be generalized in some way to other systems of music? What about pentatonic music, for example? Or the twelve-tone system? What about the richer four-note harmonies of nineteenth century music used by composers such as Chopin, Wagner, Tchaikovsky and Brahms? The answer to these questions, as we shall see, is ‘affirmative’.
Mathematicians have been studying the geometry of configurations for centuries and there is a vast literature relating to the subject. By a configuration $\{m_{r},n_{k}\}$ in $\mathbb{R}^{2}$, we mean a collection of $m$ points and $n$ lines in the Euclidean plane such that $r$ of the $n$ lines are incident with each point and $k$ of the $m$ points lie on each line. In the case of a ‘balanced’ configuration, one for which the number of points is the same as the number of lines, and where the number of points lying on each line is the same as the number of lines passing through each point, we write $\{n_{k}\}$ in place of $\{n_{k},n_{k}\}$.
@@
The triangle is a balanced configuration of type $\{3_{2}\}$ formed of three points and three lines, with two points on each line and two lines through each point. The quadrangle $\{4_{2}\}$, the pentagon $\{5_{2}\}$, the hexagon $\{6_{2}\}$ and the $n$-gon $\{n_{2}\}$ can be defined similarly.  [^2] The most well known nontrivial balanced configuration in the plane is Pappus’s $\{9_{3}\}$. By ‘nontrivial’ we mean that (unlike, say, the triangle or the quadrangle) it is not a priori obvious that it should exist. The existence of this $\{9_{3}\}$ relies on Pappus’s theorem, which states that if three points $a$, $b$, $c$ lie on a line $A$ and three more distinct points $d$, $e$, $f$ lie on a line $B$ that is distinct from $A$, then if we let $x$ denote the intersection of the lines $ae$ and $bd$, let $y$ denote the intersection of the lines $af$ and $cd$, and let $z$ denote the intersection of the lines $bf$ and $ce$, then there exists a line $C$ containing $x$, $y$, $z$. That gives nine lines altogether meeting the nine points $a,b,c,d,e,f,x,y,z$ in a $\{9_{3}\}$, which is called the Pappus configuration.   These configurations arise as the subconfigurations associated with the subgraphs of the Levi graph of the tonnetz corresponding to the musical progressions derived from hexacycles, octacycles, decacycles, dodecacycles and $2n$-cycles.
There is also the notion of a ‘complete’ quadrangle, which is a configuration that consists of four points in general position and the six lines joining them in pairs. The complete quadrangle is a configuration of type $\{4_{3},6_{2}\}$, since two points lie on each line and three lines pass through each point. By four points in ‘general position’ in $\mathbb{R}^{2}$ in we mean that no three lie on a line and no two of the lines joining them are parallel. The complete pentagon, which consists of five points in general position and the ten lines joining them in pairs, is a configuration of type $\{5_{4},10_{2}\}$. The complete hexagon is a configuration of type $\{6_{5},15_{2}\}$. These last two we meet in connection with pentatonic music and twelve-tone music. The complete $n$-gon is a configuration of type $\{n_{n-1},\mbox{$\textstyle\frac{1}{2}$}n(n-1)_{2}\}$.   [^3] We also have the dual configurations $\{6_{2},4_{3}\}$, $\{10_{2},5_{4}\}$, $\{15_{2},6_{5}\}$ obtained by interchanging the roles of points and lines. The triangle $\{3_{2}\}$ is self-dual.
@@
Our working hypothesis is that every configuration gives rise to an abstract system of music, in which there are two sets of fundamental chords (analogous to the usual major and minor triads) where certain classes of transitions between chords of one type and the other are governed by the incidence relations of the configuration. We shall be mainly concerned with balanced configurations, for which the number of chords of each type is the same, and we give examples of music systems that arise in just this way. In the case of the Eulerian tonnetz, one can represent the twelve major chords by a set of points in $\mathbb{R}^{2}$ and the twelve minor chords by a set of lines in $\mathbb{R}^{2}$ in such a way that they form a $\{12_{3}\}$. Thus, the tonnetz has the structure of an incidence geometry, and we are able to identify exactly which incidence geometry it turns out to be. One can then ask, following the program outlined above, whether other types of music can be represented as incidence geometries. We give several examples, the first being a geometry of type $\{10_{3}\}$, the so-called Desargues configuration, consisting of ten points and ten lines – in our scheme representing the ten two-note clusters and the ten three-note clusters taken from a pentatonic scale. One can think of the two-note clusters and the three-note clusters and being the major and minor chords of pentatonic music. Our second example is based on Sylvester’s theory of duads and synthemes, which can be used to construct a geometry of type $\{15_{3}\}$ formed of fifteen lines and fifteen points, known to mathematicians as the Cremona-Richmond configuration. In our scheme the underlying set from which the duads (unordered pairs) and synthemes (unordered triples of mutually exclusive duads) come is a hexachord in the sense of Schönberg, that is to say, an unordered collection of six notes taken from a twelve-tone row.  [^4] One can ask whether a syntheme – an unordered triple of unordered pairs of distinct pitches drawn from a set of six – can be meaningfully interpreted as a ‘chord’. Now, even in the common practice period, the notion of a chord admits a range of interpretation. We all agree that if several notes are played simultaneously on one instrument for the same period of time, then that constitutes a chord (possibly dissonant). But if the same notes are played as an arpeggio, as on a piano or a harp, then in many contexts this can (and indeed should) too be regarded as a chord. The arpeggiation of chords on the violin comes as a necessity, on account of the arch of the bridge; and on the guitar, chords are strummed. Even if the notes are spread over time, and over a range of voices and instruments, and in different octaves, as long as the essential unity and identity of its elements and the gesture that draws them together can be perceived, then we can call it a chord. In the case of a syntheme, we adopt a version of the same idea. If the three pairs of duads forming an entity can be recognized as a unity, the it is a syntheme, regardless of the vagaries of its presentation. Thus, the syntheme can be regarded as a natural generalization of the usual idea of a chord.
@@
We fix the chromatic scale with equal temperament and give the notes their usual names $C,C_{\sharp}/D_{\flat},D,D_{\sharp}/E_{\flat}$, etc., making no distinction between enharmonic pairs. We use curly brackets (braces) to denote an unordered set of notes. Square brackets denote an ordered set or sequence of notes. Round brackets (parentheses) are used for sequences of chords. Angled brackets are used in the case of a cycle of distinct chords, in which case the final written chord repeats the initially written chord, to signal that the sequence closes. We use the term ‘duad’ for an unordered pair of distinct notes, for example $\{C,E\}$. We write $\{C,E,G\}$ to represent an unordered triad of the notes $C$, $E$, $G$, each of which might be located within any octave. In such a context, $C$ denotes an equivalence class of notes modulo octaves. Thus, we write $CM=\{C,E,G\}$ for the ‘abstract’ $C$-major chord (i.e., the chord of pitch classes). We write $[C,E,G]$, with square brackets to denote an ordered triad $C$, $E$, $G$, with common sense specification of the octaves within which the notes appear. When we write $CM=[C,E,G]$ for an ordered triad, it is usually with a specific choice of pitch in mind for the three notes, dictated by context. For the first inversion we write $CM^{(1)}=[E,G,C]$ and for the second inversion $CM^{(2)}=[G,C,E]$. Similarly, we write $Cm=[C,E_{\flat},G]$, $Cm^{(1)}=[E_{\flat},G,C]$ and $Cm^{(2)}=[G,C,E_{\flat}]$ In a situation where the inversions are important, we write $CM^{(0)}=[C,E,G]$ for the chord in its root position. There are three further permutations of a triad, $[E,C,G]$, $[G,E,C]$ and $[C,G,E]$, making six forms altogether.  [^5] The three ‘odd’ permutations are not usually given separate names, though the perceived ‘colour’ of a chord can be sensitive to a permutation of its elements. The term ‘first inversion’ is usually applied to both of the chords with $E$ in the bass position, viz. $[E,G,C]$ and $[E,C,G]$; the term ‘second inversion’ is applied to $[G,C,E]$ and $[G,E,C]$, with $G$ in the bass. Inversions often occur in the context of four-voice part writing, where the triad, in one of its six arrangements, will have one of the voices doubled.
@@
In the case of four-note chords one has twenty-four permutations of the way in which the four notes can be distributed over the four voices. The minor sixth $Cm^{\,6}$, for example, given by $C,E_{\flat},G,A]$ in root position, admits, among its twelve odd permutations, the form $A,E_{\flat},G,C]$. The significance of this particular permutation is more apparent if we transpose the root chord down by a major third to give the $A_{\flat}m^{\,6}=[A_{\flat},C_{\flat},E_{\flat},F]$ in root position and its much-discussed odd permutation $F,C_{\flat},E_{\flat},A_{\flat}]$ which, in Wagner’s notation $F,B,E_{\flat},A_{\flat}]$ we recognize as the ‘Tristan’ chord in the form it first appears in the opening bars of the eponymous opera. The permutation is $1,2,3,4]\to[4,2,3,1]$. The same chord, at the same pitch levels, appears prominently, in the left hand, as an arpeggio, fortissimo, in a passage of Chopin’s $G$-minor Ballade, Opus 23, just before the più animato, in the enharmonically equivalent representation $E_{\sharp},B,D_{\sharp},G_{\sharp}]$, obtained by applying the same permutation to the $G_{\sharp}m^{\,6}$ chord. We shall have more to say on the Tristan chord later.  [^6] We have segued on to the consideration of minor sixths here to stress the importance of the group of permutations of $n$ objects in chord formation. For more on permutations see Callender et al (2008). We note that under the permutation $1,2,3,4]\to[4,1,2,3]$ a minor sixth becomes a half-diminished seventh. For example, $A^{\phi 7}=[A,C,E_{\flat},G]$ is the third cyclic inversion of $Cm^{6}$. Note also that under the permutation $1,2,3,4]\to[2,3,4,1]$ a minor seventh becomes a major sixth. For example, the first cyclic inversion of $Cm^{7}=[C,E_{\flat},G,B_{\flat}]$ is $E_{\flat}M^{6}=[E_{\flat},G,B_{\flat},C]$.
@@
With these conventions at hand we turn to the construction of the Eulerian tonnetz. We observe that a major chord can be changed to a minor chord in three different ways by fixing two of the tones and altering the third. For example, in the case $CM^{(0)}=[C,E,G]$, we can lower the $C$ by a semitone to obtain $Em^{(2)}=[B,E,G]$, the second inversion of the $E$-minor chord. If we lower the $E$ by a semitone we obtain $Cm^{(0)}=[C,E_{\flat},G]$, the $C$-minor chord in root position. If we raise the $G$ by a tone we obtain $Am^{(1)}=[C,E,A]$, the first inversion of $A$-minor. Similarly, a minor chord can be changed to a major chord in three different ways by fixing two of the tones and altering the third. For example, in the case $Cm^{(0)}=[C,E_{\flat},G]$, we can lower the $C$ by a tone to obtain $E_{\flat}M^{(2)}=[B_{\flat},E_{\flat},G]$ or we can raise the $E_{\flat}$ by a semitone to obtain $CM^{(0)}=[C,E,G]$ or we can raise the $G$ by a semitone to obtain $A_{\flat}M^{(1)}=[C,E_{\flat},A_{\flat}]$.
@@
In this paper we highlight – for the first time, we believe – the unifying role of Levi graphs and configurations (both geometric and combinatorial) in the analysis of music and in the creation of new forms of music. The geometry of tessellations also plays an important part. In Section II we show that the Eulerian tonnetz can be represented by a finite bipartite graph of degree three and girth six.  [^7] Here we follow the approach of Waller (1978). This work seems to have been neglected by music theorists and we have not seen Waller’s bipartite graph discussed elsewhere in the literature of the tonnetz (but see Cohn 2012 at p. 18). There is, of course, no mention of the tonnetz in Waller’s paper itself – for at that point the tonnetz had not yet become an object of renewed interest in the community of music theorists, which began with the work of Lewin and his school in the 1980s – nor does Waller mention that he is working with a Levi graph and that there should be an associated configuration. This is consistent with the fact that, with a few exceptions, such as the book of Hilbert & Cohn-Vossen (1932) and the paper of Coxeter (1948), the theory of configurations, which had been active in the nineteenth century, had been going through a dormant phase, which lasted from 1910 up to 1990 (Grünbaum 2009 at pp. 8-14). Derek Waller, a British mathematician who worked in category theory, algebra and graph theory, died in 1978 at the age of 37, so never had the opportunity to follow through on his work, as he undoubtedly would have. The connection between tonnetze and configurations is our idea, as is the realization that Waller’s construction, i.e. what we now recognize as the Eulerian tonnetz, is the Levi graph of the D222 of Daublebsky von Sterneck.  The result is stated in Proposition 1 and the graph is drawn in Figure 2. Each white vertex of this bipartite graph abuts three black vertices; each black vertex abuts three white vertices. The white vertices represent major chords; the black vertices represent minor chords. We show how properties of the four mutually disjoint hexatonic hexacycles and the three mutually disjoint octatonic octacycles can be understood in the setting of this so-called Levi graph and we comment on the role of other cycles that arise naturally in the graph, which we illustrate in Figure 3 with several examples.
@@
In Section III we show how the tonnetz can be regarded as a configuration. This is achieved by treating the transformations between major and minor chords as incidence relations in a combinatorial geometry. We reach the new and surprising conclusion that the tonnetz can be realized geometrically in the Euclidean plane $\mathbb{R}^{2}$ as a certain self-dual configuration of type $\{12_{3}\}$. Specifically, the tonnetz configuration turns out to be the $\{12_{3}\}$ of Daublebsky von Sterneck type D222. The result is stated in Proposition 2 and the configuration is drawn in Figure 4.

### v5

@@ Abstract
The Eulerian tonnetz, which is commonly represented by a network of tones, can be represented equivalently by a regular bipartite graph of degree three with twelve white vertices denoting major chords and twelve black vertices denoting minor chords. Three minor chords are associated to each major chord and three major chords are associated to each minor chord. This Levi graph can be realized geometrically as a configuration of twelve points and twelve lines in $\mathbb{R}^{2}$ with the property that three points lie on each line and three lines pass through each point. Interesting features of the tonnetz, such as the four hexatonic cycles and the three octatonic cycles, crucial for the modern understanding of nineteenth-century harmony and voice leading, can be read off directly as properties of this configuration $\{12_{3}\}$ and its Levi graph. Analogous tone networks along with their Levi graphs and configurations can be constructed for pentatonic music and twelve-tone music. These and other new tonnetze offer the promise of new methods of composition. If the constraints of the Eulerian tonnetz are relaxed to allow variations between major and minor triads at exactly two tones, the resulting bipartite graph has two components, each generating a tessellation of the plane, of a type known to Kepler, based on hexagons, squares and dodecagons. When the same combinatorial idea is applied to tetrachords of the ‘Tristan’ genus (dominant sevenths and half-diminished sevenths) the cycles of the resulting bipartite graph are sufficiently ample in girth to ensure the existence of a second configuration $\{12_{3}\}$, distinct from the Eulerian tonnetz as an incidence geometry, which can be used for a new approach to the analysis of the rich tetradic harmonies of the nineteenth century common practice.
@@ I Introduction
The well-known Eulerian tonnetz, which consists of an intricate network of relations between the twelve notes of the chromatic scale, was originally introduced by the mathematician Leonhard Euler (1739) and later developed in works of Gottfried Weber (1824), Ernst Naumann (1858), Arthur von Oettingen (1866) and Hugo Riemann (1902).  [^1] We treat the word ‘tonnetz’, spelled with a lower case ‘t’, and its plural ‘tonnetze’ as words of English.   Investigation of the tonnetz from a modern standpoint in the setting of equal temperament was carried forward in works of Waller (1978), Lewin (1982, 1987), Hyer (1989), Cohn (1996, 1997, 1998), Douthett & Steinbach (1998) and many others, attracting much attention. Figure 1 shows an alternative representation of the Eulerian tonnetz as a tessellation of the plane by regular hexagons in which the vertices are the conventional major and minor chords. This figure, viewed as an infinite graph, is familiar to music theorists – see, for example, Douthett & Steinbach (1998) at p. 248 or Tymoczko (2011b) at p. 413. But the tessellation in Figure 1 is more than a mere graph or diagram: it is a geometric object in the Euclidean plane.
@@
We propose here to introduce another geometric representation of the tonnetz – specifically, we show that the tonnetz can be visualized as a configuration of twelve lines and twelve points in $\mathbb{R}^{2}$, with the property that three of the points lie on each line and three of the lines pass through each point. The construction proceeds in several steps. First, we interpret the tonnetz as a regular bipartite graph of degree three and girth six with the two types of vertices corresponding to major and minor triads, respectively, and the edges encoding adjacency within the tonnetz. This allows for easy representation and analysis of various cyclic chord progressions that arise naturally in the tonnetz. We go on to realize this Levi graph as an incidence geometry consisting of twelve points (major triads) and twelve lines (minor triads) in the Euclidean plane, in the form of a configuration $\{12_{3}\}$, leading to interesting visualizations. Finally, we identify this configuration as the self-dual $\{12_{3}\}$ of type D222 drawn and described by Daublebsky von Sterneck (1895).
@@
The tonnetz is widely regarded as essential to the modern understanding of music based on triadic harmonies – so it is both unexpected and important to see such a connection between the tonnetz and a basic construction in Euclidean geometry. But there are challenges as well. Can the idea be generalized to other systems of music? What about pentatonic music, for example? Or the twelve-tone system? What about the richer four-note harmonies of nineteenth century common practice used in abundance by composers such as Chopin, Wagner and Tchaikovsky? The answer to these questions, as we shall see, is ‘affirmative’. Mathematicians have been studying the geometry of configurations for centuries and there is a vast literature relating to the subject. By a configuration $\{m_{r},n_{k}\}$ in $\mathbb{R}^{2}$, we mean a collection of $m$ points and $n$ lines in the Euclidean plane such that $r$ of the $n$ lines are incident with each point and $k$ of the $m$ points lie on each line. In the case of a ‘balanced’ configuration, for which the number of points is the same as the number of lines, and where the number of points lying on each line is the same as the number of lines passing through each point, we write $\{n_{k}\}$ in place of $\{n_{k},n_{k}\}$.
@@
The triangle is a balanced configuration of type $\{3_{2}\}$ formed of three points and three lines, with two points on each line and two lines through each point. The quadrangle $\{4_{2}\}$, the pentagon $\{5_{2}\}$, the hexagon $\{6_{2}\}$, and more generally the $n$-gon $\{n_{2}\}$ can be defined similarly.  [^2] The most well known nontrivial balanced configuration in the plane is Pappus’s $\{9_{3}\}$. By ‘nontrivial’ we mean that (unlike, say, the triangle or the quadrangle) it is not a priori obvious that it should exist. The existence of this $\{9_{3}\}$ relies on Pappus’s theorem, which states that if three points $a$, $b$, $c$ lie on a line $A$ and three more distinct points $d$, $e$, $f$ lie on a line $B$ that is distinct from $A$, then if we let $x$ denote the intersection of the lines $ae$ and $bd$, let $y$ denote the intersection of the lines $af$ and $cd$, and let $z$ denote the intersection of the lines $bf$ and $ce$, then there exists a line $C$ containing $x$, $y$, $z$. That gives nine lines altogether meeting the nine points $a,b,c,d,e,f,x,y,z$ in a $\{9_{3}\}$, which is called the Pappus configuration.   These configurations arise as the subconfigurations associated with the subgraphs of the Levi graph of the tonnetz corresponding to the musical progressions derived from hexacycles, octacycles, decacycles, dodecacycles, and $2n$-cycles, respectively. There is also the notion of a ‘complete quadrangle’, which is a configuration that consists of four points in general position and the six lines joining them in pairs. The complete quadrangle is a configuration of type $\{4_{3},6_{2}\}$, since two points lie on each line and three lines pass through each point. By four points in ‘general position’ in $\mathbb{R}^{2}$ we mean that no three lie on a line and no two of the lines joining them are parallel. The complete pentagon, which consists of five points in general position and the ten lines joining them in pairs, is a configuration of type $\{5_{4},10_{2}\}$. The complete hexagon is a configuration of type $\{6_{5},15_{2}\}$. These last two we meet in connection with pentatonic music and twelve-tone music. The complete $n$-gon is a configuration of type $\{n_{n-1},\mbox{$\textstyle\frac{1}{2}$}n(n-1)_{2}\}$.  [^3] We also have the dual configurations $\{6_{2},4_{3}\}$, $\{10_{2},5_{4}\}$, $\{15_{2},6_{5}\}$ obtained by interchanging the roles of points and lines. The triangle $\{3_{2}\}$ is self-dual.
@@
Our central thesis is that every configuration gives rise to an abstract system of music, in which there are two sets of fundamental tone clusters (analogous to the usual major and minor triads) where certain classes of transitions between chords of one type and the other are governed by the incidence relations of the configuration. We shall be mainly concerned with balanced configurations, for which the number of chords of each type is the same, and we give examples of music systems that arise in this way. In the case of the Eulerian tonnetz, one can represent the twelve major chords by a set of points in $\mathbb{R}^{2}$ and the twelve minor chords by a set of lines in $\mathbb{R}^{2}$ in such a way that they form a $\{12_{3}\}$. Thus, the tonnetz has the structure of an incidence geometry, and we are able to identify exactly which incidence geometry it turns out to be. One can then ask, following the program outlined above, whether other types of music can be represented as incidence geometries. We give several examples, the first being a geometry of type $\{10_{3}\}$, the Desargues configuration, consisting of ten points and ten lines – in our scheme representing the ten two-note clusters and the ten three-note clusters taken from a pentatonic scale. One can think of the two-note clusters and three-note clusters as being the ‘major’ and ‘minor’ chords of pentatonic music.  [^4] Our use of the terms ‘major’ and ‘minor’ here is for convenience, to mark the combinatorial similarity that arises naturally here to the corresponding structures in conventional Western music. As we show in Section VI, one can also consider situations for which single-note clusters are paired with multiple-note clusters. For instance, the twelve notes of the chromatic scale, regarded as single note clusters, together with the major chords, regarded as three-note clusters, form a $\{12_{3}\}$, which turns out to be a D222.
Our second example is based on Sylvester’s theory of duads and synthemes, which can be used to construct a geometry of type $\{15_{3}\}$ formed of fifteen lines and fifteen points, known as the Cremona-Richmond configuration. In our scheme the underlying set from which the duads (unordered pairs) and synthemes (unordered triples of mutually exclusive duads) come is a hexachord in the sense of Schönberg, that is to say, an unordered collection of six notes taken from a twelve-tone row.  [^5] One can ask whether a syntheme – an unordered triple of unordered pairs of distinct pitches drawn from a set of six – can be interpreted as a ‘chord’. Now, even in the common practice period, the notion of a chord admits a range of interpretation. We all agree that if several notes are played simultaneously on one instrument for the same period of time, then that constitutes a chord (possibly dissonant). But if the same notes are played as an arpeggio, as on a piano or harp, then in many contexts this can (and indeed should) too be regarded as a chord. The arpeggiation of chords on the violin comes as a necessity, on account of the arch of the bridge; and on the guitar, chords are strummed. Even if the notes are spread over time, and over a range of voices and instruments, and in different octaves, as long as the essential unity and identity of its elements and the gesture that draws them together can be perceived, then we can call it a chord. In the case of a syntheme, we adopt the same idea. If three pairs of duads forming an entity can be recognized as a unity, then it is a syntheme, regardless of the vagaries of its presentation. In this sense, the syntheme can be regarded as a natural generalization of the usual idea of a chord.
@@
We fix the chromatic scale and give the notes their usual names $C,C_{\sharp}/D_{\flat},D,D_{\sharp}/E_{\flat}$, and so on, making no distinction between enharmonic pairs. We use curly brackets (braces) to denote an unordered set of notes. Square brackets denote an ordered set or sequence of notes. Round brackets (parentheses) are used for sequences of chords. Angled brackets are used in the case of a cycle of distinct chords, in which case the final written chord repeats the initially written chord, to signal that the sequence closes. We use the term ‘duad’ for an unordered pair of distinct notes, for example $\{C,E\}$. We write $\{C,E,G\}$ to represent an unordered triad of the notes $C$, $E$, $G$, each of which might be located within any octave. In such a context, $C$ denotes an equivalence class of notes modulo octaves. Thus, we write $CM=\{C,E,G\}$ for the ‘abstract’ $C$-major chord (i.e., the chord of pitch classes).
When we write $CM=[C,E,G]$ for an ordered triad, it is usually with a specific choice of pitch in mind for the three notes, dictated by context. For the first inversion we write $CM^{(1)}=[E,G,C]$ and for the second inversion $CM^{(2)}=[G,C,E]$. Similarly, we write $Cm=[C,E_{\flat},G]$, $Cm^{(1)}=[E_{\flat},G,C]$ and $Cm^{(2)}=[G,C,E_{\flat}]$. In a situation where the inversions are important, we write $CM^{(0)}=[C,E,G]$ for the chord in its root position. There are three further permutations of a triad, $[E,C,G]$, $[G,E,C]$ and $[C,G,E]$, making six forms altogether.  [^6] The three ‘odd’ permutations are not usually given separate names, though the perceived ‘colour’ of a chord can be sensitive to a permutation of its elements. The term ‘first inversion’ is usually applied to both of the chords with $E$ in the bass position, viz. $[E,G,C]$ and $[E,C,G]$; the term ‘second inversion’ is applied to $[G,C,E]$ and $[G,E,C]$, with $G$ in the bass. Inversions often occur in the context of four-voice part writing, where the triad, in one of its six arrangements, will have one of the voices doubled.
@@
In the case of four-note chords one has twenty-four permutations of the way in which the four notes can be distributed over the four voices. The minor sixth $Cm^{\,6}$, for example, given by $C,E_{\flat},G,A]$ in root position, admits, among its twelve odd permutations, the form $A,E_{\flat},G,C]$. The significance of this permutation is more apparent if we transpose the root chord down by a major third to give the $A_{\flat}m^{\,6}=[A_{\flat},C_{\flat},E_{\flat},F]$ in root position and its much-discussed odd permutation $F,C_{\flat},E_{\flat},A_{\flat}]$ which, in Wagner’s notation $F,B,E_{\flat},A_{\flat}]$ we recognize as the ‘Tristan’ chord in the form it first appears in the opening bars of the eponymous opera. The permutation is $1,2,3,4]\to[4,2,3,1]$. The same chord, at the same pitch levels, appears prominently, in the left hand, as an arpeggio, fortissimo, in a passage of Chopin’s $G$-minor Ballade, Opus 23, just before the più animato, in the enharmonically equivalent representation $E_{\sharp},B,D_{\sharp},G_{\sharp}]$, obtained by applying the same permutation to the $G_{\sharp}m^{\,6}$ chord.
We have segued on to the consideration of minor sixths here to stress the importance of the group of permutations of $n$ objects in chord formation.   [^7] See Callender et al (2008). We note that under the permutation $1,2,3,4]\to[4,1,2,3]$ a minor sixth becomes a half-diminished seventh. For example, the $Am$ half-diminished seventh $A^{\phi 7}=[A,C,E_{\flat},G]$ is the third cyclic inversion of $Cm^{6}$. Note also that under the permutation $1,2,3,4]\to[2,3,4,1]$ a minor seventh becomes a major sixth. For example, the first cyclic inversion of $Cm^{7}=[C,E_{\flat},G,B_{\flat}]$ is $E_{\flat}M^{6}=[E_{\flat},G,B_{\flat},C]$.  We shall have more to say on the Tristan chord later.
@@
With these conventions at hand we turn to the construction of the Eulerian tonnetz. We observe that a major chord can be changed to a minor chord in three different ways by fixing two of the tones and altering the third. For example, in the case $CM^{(0)}=[C,E,G]$, we can lower the $C$ by a semitone to obtain $Em^{(2)}=[B,E,G]$. If we lower the $E$ by a semitone we obtain $Cm^{(0)}=[C,E_{\flat},G]$. If we raise the $G$ by a tone we obtain $Am^{(1)}=[C,E,A]$.
Similarly, a minor chord can be changed to a major chord in three different ways by fixing two of the tones and altering the third. For example, in the case of the chord $Cm^{(0)}=[C,E_{\flat},G]$, we can lower the $C$ by a tone to obtain $E_{\flat}M^{(2)}=[B_{\flat},E_{\flat},G]$ or we can raise the $E_{\flat}$ by a semitone to obtain $CM^{(0)}=[C,E,G]$ or we can raise the $G$ by a semitone to obtain $A_{\flat}M^{(1)}=[C,E_{\flat},A_{\flat}]$.
@@
In this paper we highlight – for the first time, we believe – the unifying role of Levi graphs and configurations (both geometric and combinatorial) in the analysis of music and in the creation of new forms of music. The geometry of tessellations also plays an important part. In Section II we show that the Eulerian tonnetz can be represented by a finite bipartite graph of degree three and girth six.  [^8] Here we follow the approach of Waller (1978). This work seems to have been neglected by music theorists and we have not seen Waller’s bipartite graph discussed elsewhere in the literature of the tonnetz (but see Cohn 2012 at p. 18). There is, of course, no mention of the tonnetz in Waller’s paper itself – for at that point the tonnetz had not become an object of renewed interest in the community of music theorists, which began with the work of Lewin and his school in the 1980s – nor does Waller mention that he is working with a Levi graph and that there should be an associated configuration. This is consistent with the fact that, with a few exceptions, such as the book of Hilbert & Cohn-Vossen (1932) and the papers of Coxeter (1950, 1958), the theory of configurations, which had been active in the nineteenth century, had been going through a dormant phase, lasting from 1910 up to 1990 (Grünbaum 2009 at pp. 8-14). Derek Waller, a British mathematician who worked in category theory, algebra and graph theory, died in 1978 at the age of 37. The connection between tonnetze and configurations is our idea, as is the realization that Waller’s construction, i.e. what we now recognize as the Eulerian tonnetz, is the Levi graph of the D222 of Daublebsky von Sterneck.  The result is stated in Proposition 1 and the graph is drawn in Figure 2. Each white vertex of this bipartite graph abuts three black vertices; each black vertex abuts three white vertices. The white vertices represent major chords; the black vertices represent minor chords. We show how properties of the four mutually disjoint hexatonic cycles and the three mutually disjoint octatonic cycles can be understood in the setting of this Levi graph and we comment on the role of other cycles that arise naturally in the graph, which we illustrate in Figure 3 with several examples.
@@
In Section III we show how the tonnetz can be regarded as a configuration. This is achieved by treating the transformations between major and minor chords as incidence relations in a combinatorial geometry. We reach the surprising conclusion that the tonnetz can be realized geometrically in the Euclidean plane $\mathbb{R}^{2}$ as a certain self-dual configuration of type $\{12_{3}\}$. Specifically, the tonnetz turns out to be the configuration $\{12_{3}\}$ of Daublebsky von Sterneck type D222. The result is stated in Proposition 2 and the configuration of major and minor chords is drawn in Figure 4.